Which would have the greatest effect on the money supply: 100 deposited when the required

reserve ratio is 10% or 80 deposited when the ratio is 8%? If you can explain it with the money multiplier formula of C+D/C+ER+RR that would be great if possible.  
Answer #1

When the required reserve ratio is 10%, simple deposit multiplier is 1/10% = 10. Now increase in deposits is 100. Hence final increase in money supply is 100*10 = 1000

When the required reserve ratio is 8%, simple deposit multiplier is 1/8% = 12.50. Now increase in deposits is 80. Hence final increase in money supply is 80*12.50 = 1000

Therefore increase in money supply is same under both conditions. This happens because we are using only simple deposit multiplier, since we do not have information about excess reserves of the banking system.